OpenStudy (anonymous):

G is jointly proportional to the square roots of m, n, and p. If m = 1, n = 9, and p = 16, then G = 72 find the constant of prportionality

OpenStudy (mathmath333):

\(\large\tt \color{black}{G=k(\sqrt m \sqrt n \sqrt p)}\) \(\large\tt \color{black}{72=k(\sqrt1\sqrt9\sqrt16)}\) \(\large\tt \color{black}{k=\dfrac{72}{(\sqrt1\sqrt9\sqrt{16})}}\)